OpenStudy (anonymous):

Write the standard equation for each Hyperbola. 1. x2 - 100y2 = 100 2. 4y2 - 25x2 = 100 3. vertices (-6, 0) ( 6, 0) and focus (-8, 0) ( 8, 0) 4. vertices ( 0, -12) ( 0, 12) and co-vertices (-11, 0) (11, 0)

OpenStudy (anonymous):

Write the standard equation of each Ellipse. Find the center, vertices, co-vertices and foci. 1. 25x2 + 9y2 = 225 2. 49x2 + y2 = 49 3. x2 - 2x + 9y2 - 8 = 0 4. x2 + 4y2 - 18x -8y + 81 = 0

OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros. 1) -2, -1, 1 2) i, 4 3) i, 2 - √3 4) 1, 4, 1 + √2

OpenStudy (anonymous):

f(x) = anxn + an - 1...+ a1 + a0: If n is odd and an > o, then f(x) approaches _____as x approaches positive infinity and f(x) approaches _____ as x approaches negative infinity. If n is odd and an < 0, then f(x) approaches _____ as x approaches positive infinity and f(x) approaches _____ as x approaches negative infinity. If n is even and an > 0, then f(x) approaches _____ as x approaches positive infinity and f(x) approaches _____ as x approaches negative infinity. If n is even and an < 0, then f(x) approaches _____ as x approaches positive infinity and f(x) approaches _____ as x approaches negative infinity.

OpenStudy (anonymous):

Somebody please help !

OpenStudy (mertsj):

What test is this?

OpenStudy (anonymous):

Its not a test it my home work and im lost @Mertsj

OpenStudy (mertsj):

It has been my experience that homework does not typically address such a wide variety of topics.